OpenStudy (anonymous):

Write y = x + 5 in standard form using integers. A. –x – 6y = 30 B. 6x – y = 30 C. –x + 6y = 30 D. –x + 6y = 5

OpenStudy (anonymous):

@.Sam. help please

sam (.sam.):

I think your question has a mistake

OpenStudy (anonymous):

really?

sam (.sam.):

yes

OpenStudy (anonymous):

what should I do it's an exam

sam (.sam.):

Did you write your question correctly?

OpenStudy (anonymous):

is there an answer close I can't its online

sam (.sam.):

How about you draw it out?

OpenStudy (texaschic101):

standard form : Ax + By = C y = x + 5 -x + y = 5 x - y = 5 I am going to go with B. You might need to get a second opinion.